OpenStudy (agl202):

Medal Fan for best response!!! Line AB has an equation of a line y = 5x - 2, and line DC has an equation of a line y = -5x - 2. These two lines are... parallel because the slopes of the lines are the same perpendicular because the product of the slopes is -1 both representations of the same line neither parallel nor perpendicular

OpenStudy (trojanpoem):

y -5x + 2 = 0 Slope = 5/1 = 5 y +5 x + 2 = 0 Slope = -5/1 = -5 neither parallel nor perpendicular

OpenStudy (anonymous):

i think is d because they are not parallel nor perpendicular

OpenStudy (agl202):

U guys sure?

OpenStudy (trojanpoem):

I am sure.

OpenStudy (trojanpoem):

You have two lines if the product of both slopes is -1 then perp if slope 1 = slope 2 parrallel

OpenStudy (holly00d1248):

I'd agree with them if i were you

OpenStudy (trojanpoem):

If you don't understand, ask for explanition

OpenStudy (agl202):

Okay...Thx : )!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

I think this is pretty accurate :P